OpenStudy (anonymous):

How do I solve sin 2x = 10/4?

OpenStudy (anonymous):

sin 2x = 10/4 sin 2x = 2.5 \[2x = \sin^-1(2.5)\]

OpenStudy (dape):

You use the inverse function of sin, the arcsin function, it has the property \[ \arcsin(\sin(x))=x \] So that \[ \arcsin(\sin(2x))=2x=\arcsin\left(\frac{10}{4}\right) \\ x=\frac{1}{2}\arcsin(2.5) \]